# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

### v1

@@ V-D3 Reward Model
  • $\text{RUL}_{\text{gain}}(S_{t},a_{t})$ represents the gain in RUL as a consequence of taking action $a_{t}$ at state $S_{t}$. This gain is computed using the state-transition probabilities $p(S_{t+1}|S_{t},a_{t})$ derived from the C-MAPSS dataset. A higher RUL gain is desirable as it suggests an extension in the machine’s operational lifespan.
@@
  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on real-world operation costs and directly affects the overall expenditure.
@@
  • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$. Downtime is critical in operational settings where machine availability significantly impacts productivity or service levels.
@@
  • $\alpha,\beta,\gamma$ are hyperparameters that weigh the importance of RUL, cost, and downtime, respectively. By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL of machinery, minimizing costs, or reducing downtime, as per the organizational objectives.
V-D4 Hyperparamters
To train the algorithm, we use several hyperparameters that were carefully chosen based on empirical experiments. In the following, we present detailed hyperparameters for each DRL algorithm.
@@ V-D4 Hyperparamters
Deep Q-network (DQN). DQN alleviates the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function. It also introduces experience replay to improve learning stability and efficiency. The learning rate was set to $0.0001$, the discount factor was set to $0.99$, and the replay buffer size was set to $1,000,000$ experiences. We used a batch size of $512$ and updated the target network every $50$ steps. The exploration vs. exploitation trade-off was controlled by an epsilon value that started at $1.0$ and decayed over time to a final value of $0.01$. Additionally, we used a neural network with two hidden layers of $256$ units each, and we used the Adam optimizer with a learning rate of $0.0001$.
@@
Proximal Policy Optimization (PPO). PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy. We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$. A clip range of $0.2$ was used to limit the updates to the policy, and we applied a value function coefficient of $0.5$ to balance the importance of the policy and the value function. Additionally, we set the entropy coefficient to $0.01$ to encourage exploration, and we used a batch size of $64$. We optimize the policy every 10 epochs and used the Adam optimizer with an epsilon value of $0.00001$. Finally, we set the GAE parameter to $0.95$ to estimate the advantage function.
@@
Soft Actor-Critic (SAC). SAC is particularly effective for problems with high-dimensional action spaces. SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient. The learning rate for the actor and critic networks was set to $0.0003$, and we used a discount factor of $0.99$. We applied a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function. The target entropy was set to $-2.0$, and we used a replay buffer of $1,000,000$ experiences with a batch size of $256$. We also applied a soft update of the target networks with a rate of $0.005$, and we initialized the log standard deviation of the policy with a value of $-2$. We trained the algorithm for $100,000$ time steps and used the Adam optimizer with an epsilon value of $0.0001$.
V-E Results and Discussion
We compare the performance of DQN [22], PPO [23], and SAC [24] on our designed task in the Figure 5, Figure 6, and Figure 7. The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes. The experimental results show that PPO outperforms both DQN and SAC, achieving higher total rewards on the task. The average total rewards obtained by PPO are $55,000$, while DQN and SAC achieve average total rewards of $50,000$ and $45,000$, respectively. However, DQN converges faster than PPO and SAC.
In general, PPO and SAC are considered to be more sample-efficient than DQN, which means that they can achieve comparable performance with fewer samples. This is because PPO and SAC use more efficient updates that make better use of the available data. In contrast, DQN can be more sensitive to the choice of hyperparameters and the amount of data used during training, which can affect its convergence speed.
However, the performance of these algorithms can vary depending on the specific task and the quality of the implementation. In some cases, DQN may converge faster than PPO or SAC, especially if the action space is discrete and the observations are low-dimensional. On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster. Results that are shown in Figure 5 and Figure 7 have confirmed this theory.
@@ V-E Results and Discussion
Figure 7: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in Cartpole environment.
@@ VI Conclusions
In this paper, we have presented a comprehensive prescriptive maintenance framework that uniquely integrates real-time RUL estimations with reinforcement learning algorithms for dynamic action recommendations. The framework starts by employing the transformer algorithm to estimate the RUL distribution of engines, offering a robust, adaptive approach to understanding machine health in real-time. As machinery operates, our model continues to dynamically update these RUL estimates, providing an ever-evolving understanding of the equipment’s condition.
We rigorously compared the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates. Our findings indicate that the transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process.
For recommending maintenance actions, we used SAC, DQN, and PPO. Each of these algorithms has its strengths and weaknesses, but they all were adapted to work seamlessly with the dynamically updated RUL estimates. Our experimental results showed that SAC excelled in scenarios where the cost function was highly non-linear, while PPO and DQN were more suited for environments with more predictable dynamics.
This framework opens up new avenues for research, particularly in the fine-tuning of hyperparameters that govern the trade-offs between extending the machinery’s useful life and minimizing operational costs. Future work could involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machinery.
In conclusion, our work signifies a step forward in the quest for more intelligent, adaptive, and cost-effective prescriptive maintenance systems. By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a holistic solution that dynamically adapts to the operational realities, thereby driving both economic and operational efficiencies.

### v2

@@ V-C3 Reward Model
  • $\text{RUL}_{\text{gain}}(S_{t},a_{t})$ represents the gain in RUL as a consequence of taking action $a_{t}$ at state $S_{t}$. This gain is computed using the state-transition probabilities $p(S_{t+1}|S_{t},a_{t})$ derived from the C-MAPSS dataset. A higher RUL gain is desirable as it suggests an extension in the machine’s normal working time.
@@
  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on experiences.
@@
  • $\text{DownTime}(a_{t})$ signifies the machine downtime resulting from action $a_{t}$. Downtime is critical in operational settings, which impacts productivity or service levels.
@@
  • $\alpha,\beta,\text{and}~{}\gamma$ are hyperparameters that weigh the importance of RUL, cost, and downtime, respectively. By fine-tuning these hyperparameters, the model can be customized to focus on extending the RUL, minimizing costs, or reducing downtime.
V-C4 Hyperparamters
For each DRL algorithm embedded in Algorithm 4, we carefully choose its corresponding hyperparameters based on empirical experiments. Details for each DRL algorithm are given as follows.
@@ V-C4 Hyperparamters
Deep Q-network (DQN). DQN reduces the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function [35]. It also introduces experience replay to improve learning stability and efficiency. The learning rate is set to $10^{-4}$, the discount factor is set to $0.99$, and the replay buffer size is set to $10^{6}$ experiences. We use a batch size of $512$ and update the target network every $50$ steps. The exploration-exploitation trade-off is controlled by an $\epsilon$ value that started at $1.0$ and decayed over time to a final value of $0.01$. Additionally, we use a neural network with two hidden layers of $256$ units each, and we utilize the Adam optimizer with a learning rate of $10^{-4}$.
@@
Proximal Policy Optimization (PPO). PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy [36]. We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$. A clip range of $0.2$ is used to limit the updates to the policy, and we apply a value function coefficient of $0.5$ to balance the importance of the policy and the value function. Additionally, we set the entropy coefficient to $0.01$ to encourage exploration and a batch size as $64$. Then, we optimize the policy every 10 epochs and use the Adam optimizer with a learning rate of $0.0001$. Finally, GAE parameter is set to $0.95$ to estimate the advantage function.
@@
Soft Actor-Critic (SAC). SAC is particularly effective for problems with high-dimensional action spaces [37]. SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient. The learning rate for the actor and critic networks is set to $0.0003$, and we use a discount factor of $0.99$. Then, we apply a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function. The target entropy is set to $-2.0$, and we use a replay buffer of $10^{6}$ experiences with a batch size of $256$. We also apply a soft update of the target networks with a rate of $0.005$, and we initialize the log standard deviation of the policy with a value of $-2$. Finally, we train the algorithm for $10^{5}$ time steps and use the Adam optimizer with a learning rate of $0.0001$.
V-D Results and Discussion
@@
We compare the performance of DQN, PPO, and SAC on our designed task in Figures 1213. The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes. The experimental results show that PPO outperforms both DQN and SAC, achieving higher total rewards on the task. The average total rewards obtained by PPO are $55,000$, while DQN and SAC achieve average total rewards of $50,000$ and $45,000$, respectively. However, DQN converges faster than PPO and SAC.
In general, PPO and SAC are considered to be more sample-efficient than DQN, which means that they can achieve comparable performance with fewer samples. This is because PPO and SAC use more efficient updates that make better use of the available data. In contrast, DQN can be more sensitive to the choice of hyperparameters and the amount of data used during training, which can affect its convergence speed. However, the performance of these algorithms can vary depending on the specific task and the quality of the implementation. In some cases, DQN may converge faster than PPO or SAC, especially if the action space is discrete and the observations are low-dimensional. On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster.
Figure 14: Differences between RL and RLHF.
V-E Reinforcement Learning with Human Feedback
Incorporating human expertise into the reinforcement learning process helps to resolve some of the challenges encountered with traditional RL methods. Especially, in scenarios where real-world rewards are hard to quantify or can be misleading, human feedback provides a way to guide the learning process more effectively. The comparison between RL and RLHF is presented in Fig. 14.
As machines and their associated systems become increasingly complex, the conditions that need to be considered maybe take a long time for Algorithm 5. Human experts, with their working experiences and domain knowledge to find complex patterns, can provide valuable insights that may not be readily available through raw data. By integrating human feedback into the RL process, the framework can:
  • Augment the learning process with domain-specific knowledge.
  • Address the cold start problem where initial exploration might be risky or expensive.
  • Improve learning efficiency by reducing the exploration of sub-optimal actions.
Thus, we further incorporate the following information into our proposed framework in Algorithm 5:
  1. 1.
    Human Opinions: Collect human predictions of the desired behavior. Experienced engineers can provide opinions on maintenance decisions based on various machine states from the C-MAPSS dataset.
  2. 2.
    Feedback on Agent’s Actions: After the agent makes a decision, a human can provide feedback on the chosen action, guiding the agent on the appropriateness of its actions.
Then, we introduce a term in the reward function that captures human feedback
$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if human feedback is positive},\\ r_{\text{negative}},&\text{if human feedback is negative},\\ 0,&\text{otherwise},\end{cases}$ (7)
based on which, the modified reward function for DRL is
$R(S_{t},a_{t},S_{t+1})$ $=\alpha\times\text{RUL}_{\text{gain}}(S_{t},a_{t})-\beta\times\text{Cost}(a_{t})-$
$\gamma\times\text{DownTime}(a_{t})+\delta\times R_{\text{human}}(a_{t}),$
where $\delta$ is a hyperparameter weighing the importance of human feedback.
Input: Observation $\bm{s}$, Exploration rate $\epsilon$, Learning rate $\alpha$, Discount factor $\gamma$, Human feedback 00.
Output: Optimal action $a$ for maintenance
1 Initialize the neural network weights;
2 Initialize a mechanism to incorporate human feedback (e.g., reward shaping, action recommendation);
3 for each episode do
       /* Incorporate Human Demonstrations */
4       If available, utilize human demonstrations to set initial policy;
5       Initialize the starting state $\bm{s}$;
6       while $s$ is not a terminal state do
             /* Action Selection with Human Feedback */
7             Select action $a$ using an $\epsilon$-greedy policy or based on human recommendations;
8             Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
             /* Adjust Reward Based on Human Feedback */
9             Incorporate human feedback into the reward:
$R_{\text{human}}(a_{t})=\begin{cases}r_{\text{positive}},&\text{if HF is positive},\\ r_{\text{negative}},&\text{if HF is negative},\\ 0,&\text{otherwise}.\end{cases}$
Adjust reward $r$ based on $R_{\text{human}}(a_{t})$;
             /* Network Update */
10             Calculate TD error $\delta=r+\gamma\max_{a^{\prime}}Q(\bm{s^{\prime}},a^{\prime})-Q(\bm{s},a)$;
11             Update neural network weights using $\delta$ and learning rate $\alpha$;
             /* Adjustment of Exploration and State */
12             Adjust the exploration rate $\epsilon$;
13             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
14
15       end while
      /* Incorporate Feedback into Learning Process */
16       Periodically adjust the learning process based on accumulated human feedback;
17
18 end for
return Optimal maintenance action $a$.
Algorithm 5 The RLHF Algorithm for Maintenance Actions Selection
V-F Results for RLHF
@@
Figure 15: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
Figure 16: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
Figure 15 and Figure 16 reflect the changes after adding on the human feedback in the reward function. The results show that the speed of convergence does not improve too much, but the process becomes more stable and variances minimized. This is affected by the weight of human feedback in the reward function as well as the complexity of the task. Since our task converges very fast even without human feedback, there is not much room for improvement. In the future, we will conduct experiments on more complex real-world scenarios to confirm the effectiveness of the RLHF algorithm.
@@ VI Conclusions
In this paper, we present a comprehensive prescriptive maintenance framework that integrates real-time RUL estimations with DRL algorithms for dynamic action recommendations. The framework starts by employing the Transformer network to estimate the RUL distribution of engines, offering a robust and adaptive data-driven approach to understanding machine conditions in real-time. As machine operates, our model continues to dynamically updating RUL estimates, providing an ever-evolving status of the equipment’s condition.
We rigorously compare the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates. Our findings indicate that the Transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process. For DRL-based maintenance action recommendation, we demonstrate the efficiency of SAC, DQN, and PPO. Each of these algorithms has its strengths and weaknesses, but they all are adapted to working seamlessly with the dynamically updated RUL estimates. Future work may involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machines.
In conclusion, our work signifies a step forward in the request for more intelligent, adaptive, and cost-effective prescriptive maintenance systems. By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a solution that dynamically adapts to the operational realities, thereby enhancing both economic and operational efficiencies.